Find the volume of a refrigerator that is a rectangular solid and is 2 meters tall and has a base that is 1.2 meters square.
A) 1.44 cubic meters B) 2.4 cubic meters C) 2.88 cubic meters D) 3.2 cubic meters

Answer by jsmallt9(2063)

Volume (V) of a rectangular solid (prism) is either l*w*h or B*h where
l = length
w = width
h = height
B = area of the base

We can use either formula. The only tricky part of this problem is in interpreting "a base that is 1.2 meters square". This does not mean that the base has an area that is 1.2 square meters. It means that the base is a square with sides of 1.2 meters! (IOW: "meters square" is not the same as "square meters".)

Now we can solve the problem:
V = l*w*h
l = 1.2
w = 1.2
h = 2
V = (1.2)*(1.2)*(2) = 2.88 cu. meters

If we want to use the second formula instead we need to find the area of the base (B). Since the base is a square the area is:
B = (1.2)*(1.2) = 1.44
Then
V = B*h
V = (1.44)*(2) = 2.88 cu. meters